use the square root property to solve the equation

(1) (x+1)^2=9

(2) x^2=121

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

problem 1:
(x+1)^2 = 9
take the square root of both sides of this equation to get:
x+1 = 3
subtract 1 from both sides of this equation to get:
x = 2
problem 2:
x^2 = 121
take the square root of both sides of this equation to get:
x = 11

Answer by ikleyn(53942)

.


        In his post, tutor @Theo lost some solutions to given equations.
        So,  his treatment of the problem is incomplete.
        I came to bring a correct complete solutions.


- - - - - - - - Equation (1) - - - - - - - -

(x+1)^2 = 9


Take square root of both sides.  You will get

(x+1) = +/-3.


It means that we have two cases.


Case 1.  x+1 = 3.  Then  x = 3-1 = 2.

Case 2.  x+1 = -3.  Then x = -3-1 = -4.


So, equation (1) has two solutions.  They are  x = 2  and  x = -4.


Check.  Substitute these values into equation (1) to make sure that they suit perfectly.



- - - - - - - - Equation (2) - - - - - - - -

x^2 = 121


Take square root of both sides.  You will get

x = +/-11.


It means that the equation has two solutions,  x = 11 and x = -11.


CHECK. Substitute these values into equation (2) to make sure that they suit perfectly.

Solved correctly.

------------------------------------

It might seem the loss of roots is a minor issue.

                In fact,  it is not so.

Loss of roots is a failed test of understanding square roots.
So, solving this equation correctly or incorrectly is an easy way
to check if a person does understand the subject.
In 10 seconds, everything became clear.